OpenStudy (anonymous):

A line passes through the origin and has a slope of –3. What is the equation of the line that is perpendicular to the first line and passes through the point (3, 4)?

OpenStudy (anonymous):

The slope of the line is the opposite recipricol of the first line: 1/3 now you have y=1/3x+b so you know the slope now just plug in x,y coordinate to the slope intercept form to find b 4=1/3(3)+b 4=1+b b=3 So the equation is y=1/3x+3